You can multiply two numbers by rewriting the numbers as the difference of two squares. For example: 36 x 24 = (30 + 6)(30 - 6) = 30^2 - 6^2 = 900 - 36 = 864
Use this method to multiply 27 x 19. Explain how you rewrote the numbers..

Answer by mananth(16949)

Use this method to multiply 27 x 19. Explain how you rewrote the numbers..
27+19=46
46/2 =23
...
(23+4)(23-4)
=23^2-4^2
=529-16
=513